Find the conical form of a quadratic form Q(x,y)=2x^2+2y^2-2xy by using an orthogonal transformation hence find nature,rank,index and signature of the conical form?

1
Expert's answer
2022-02-08T12:32:19-0500

"\displaystyle
Q(x,y)=2x^2-2xy+2y^2=2x^2-xy-yx+2y^2\\"

"\displaystyle
=\begin{pmatrix}
 x & y 
\end{pmatrix}\begin{pmatrix}
 2 & -1 \\
 -1 & 2
\end{pmatrix}\begin{pmatrix}
 x \\
 y
\end{pmatrix}"

Since "\displaystyle
A=\begin{pmatrix}
 2 & -1 \\
 -1 & 2
\end{pmatrix}" is symmetric then it "A" is orthogonally diagonalizable and has real eigenvalues.

Now, the eigenvalues of "A" are "\displaystyle
\lambda_1=1\text{ and }\lambda_2=3."


An eigenvector corresponding to the eigenvalue "\displaystyle
\lambda_1=1" is;

"\displaystyle
\begin{pmatrix}
 1 \\
 1
\end{pmatrix}"

and an eigenvector corresponding to the eigenvalue "\displaystyle
\lambda_2=3" is;

"\displaystyle
\begin{pmatrix}
 1 \\
 -1
\end{pmatrix}"

Hence, the eigenvectors "\displaystyle
(1,\ 1)\text{ and }(1,\ -1)" form an orthogonal basis for "\R^2". Normalizing each of these eigenvectors produces an orthonormal basis;

"\displaystyle
p_1=\frac{(1,\ 1)}{\|(1,\ 1)\|}=\left(\frac{1}{\sqrt{2}},\ \frac{1}{\sqrt{2}}\right)", "\displaystyle
p_2=\frac{(1,\ -1)}{\|(1,\ -1)\|}=\left(\frac{1}{\sqrt{2}},\ -\frac{1}{\sqrt{2}}\right)"


Using "\displaystyle
p_1 \text{ and }p_2" as column vectors, construct the matrix "P" such that;

"\displaystyle
P=\begin{pmatrix}
 \frac{1}{\sqrt{2}} & \frac{1}{\sqrt{2}} \\
 \frac{1}{\sqrt{2}} & -\frac{1}{\sqrt{2}}
\end{pmatrix}"

Hence, the transformation "\displaystyle
\begin{pmatrix}
 x \\
 y
\end{pmatrix}=\begin{pmatrix}
 \frac{1}{\sqrt{2}} & \frac{1}{\sqrt{2}} \\
 \frac{1}{\sqrt{2}} & -\frac{1}{\sqrt{2}}
\end{pmatrix}\begin{pmatrix}
 m \\
 n
\end{pmatrix}" transforms "Q(x,y)" to sum of squares or canonical form.


That is;

"\displaystyle
Q(x,y)=\begin{pmatrix}
 x & y 
\end{pmatrix}\begin{pmatrix}
 2 & -1 \\
 -1 & 2
\end{pmatrix}\begin{pmatrix}
 x \\
 y
\end{pmatrix}"

"\displaystyle
\Rightarrow Q(m,n)=\left(\begin{pmatrix}
 \frac{1}{\sqrt{2}} & \frac{1}{\sqrt{2}} \\
 \frac{1}{\sqrt{2}} & -\frac{1}{\sqrt{2}}
\end{pmatrix}\begin{pmatrix}
 m \\
 n
\end{pmatrix}\right)^T\begin{pmatrix}
 2 & -1 \\
 -1 & 2
\end{pmatrix}\begin{pmatrix}
 \frac{1}{\sqrt{2}} & \frac{1}{\sqrt{2}} \\
 \frac{1}{\sqrt{2}} & -\frac{1}{\sqrt{2}}
\end{pmatrix}\begin{pmatrix}
 m \\
 n
\end{pmatrix}"

"\displaystyle
=(m\ \ n)\begin{pmatrix}
 \frac{1}{\sqrt{2}} & \frac{1}{\sqrt{2}} \\
 \frac{1}{\sqrt{2}} & -\frac{1}{\sqrt{2}}
\end{pmatrix}^T\begin{pmatrix}
 2 & -1 \\
 -1 & 2
\end{pmatrix}\begin{pmatrix}
 \frac{1}{\sqrt{2}} & \frac{1}{\sqrt{2}} \\
 \frac{1}{\sqrt{2}} & -\frac{1}{\sqrt{2}}
\end{pmatrix}\begin{pmatrix}
 m \\
 n
\end{pmatrix}"

"\displaystyle
=(m\ \ n)\begin{pmatrix}
 1 & 0 \\
 0 & 3
\end{pmatrix}\begin{pmatrix}
 m \\
 n
\end{pmatrix}""\displaystyle
=m^2+3n^2"

which is the canonical form of the given quadratic form.


Nature: it is positive definite in nature since all eigenvalues are positive numbers.


Rank: of a quadratic from is the number of square terms in its canonical form. So rank of "Q(x,y)=2"


Index: of a quadratic from is the number of positive square terms in its canonical from. So the index of "Q(x,y)=2"

Signature: of "Q(x,y)=" (number of positive square terms)- (number of negative square terms)"=2"